OpenStudy (anonymous):

Solve x2 + 7x + 2 = 0.

OpenStudy (anonymous):

do you mean \[x^{2} +7x +2 = 0\]

OpenStudy (mimi_x3):

\[x = \frac{-(7)\pm\sqrt{(7)^{2}-4(1)(2)}}{2(1)} = \frac{-7\pm\sqrt{41}}{2} \]

OpenStudy (anonymous):

yes i do!